OpenStudy (anonymous):

L'Hopital's Rule: lim x-> -inf of (arctan x) / (2^x)? Please help....:'(

OpenStudy (anonymous):

pretty sure you do not use l'hopital for this

OpenStudy (anonymous):

Really?....but this section is on L'Hopital's Rule..

OpenStudy (anonymous):

\[\lim_{x\to -\infty}\tan^{-1}(x)=-\frac{\pi}{2}\] so it is not in the form \(\frac{\infty}{\infty}\) it may be in that section, they are checking to see if you know when you can and cannot use l'hopital

OpenStudy (anonymous):

Oh... you are right. Thanks!

OpenStudy (anonymous):

you can only use it in the form \(\frac{0}{0}\) or \(\frac{\infty}{\infty}\) if you try to use it here you will get the wrong answer

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

We have 0 * infinity too for L'Hopital's, but I didn't think that would be it for this problem.. thank you!

OpenStudy (anonymous):

well not really if it looks like \(0\times \infty\) to have to change it to make it look like \(\frac{0}{0}\) or \(\frac{\infty}{\infty}\)